# Difference-in-Differences Designs: A Practitioner’s Guide

### v1

@@ 3.2 Identifying assumptions: parallel trends
A research design is a strategy—a set of assumptions—to identify and estimate specific target parameters. Many different assumptions can identify the missing counterfactual for $ATT(2014)$ in the Medicaid example. For example, mean independence between $Y_{i,2014}(0)$ and $D_{i}$ implies that the counterfactual equals average 2014 mortality rates in non-expansion counties ($\mathop{}\!\mathbb{E}_{\omega}[Y_{i,2014}(0)|D_{i}=0]$). Under this assumption, which essentially entails assuming that Medicaid expansion is as-good-as random, the cross-sectional mortality gap in 2014 between expansion and non-expansion counties is the $ATT(2014)$. Similarly, time invariance of $Y_{i,t}(0)$ among expansion counties (plus the fact that we ruled out anticipatory behavior) implies that the counterfactual equals 2013 mortality rates in expansion counties ($\mathop{}\!\mathbb{E}_{\omega}[Y_{i,2013}(0)|D_{i}=1]$). Under this assumption, which essentially rules out non-treatment-related changes in the outcome variable, the “time trend” in average mortality in expansion counties is the $ATT(2014)$.
DiD comes from an alternative assumption that identifies the relevant counterfactual even when the mean of $Y_{i,t=2}(0)$ differs across treatment groups (which violates mean independence) and changes over time (which violates time invariance). The so-called parallel trends assumption states that, in the absence of treatment, the average outcome evolution is the same among treated and comparison groups.
@@
The (weighted) average change of $Y_{i,t=2}(0)$ from $Y_{i,t=1}(0)$ is the same between treated and comparison groups, i.e.,
@@
If parallel trends holds, then it is easy to construct $\mathop{}\!\mathbb{E}_{\omega}[Y_{i,t=2}(0)|D_{i}=1]$ from observable quantities—that is, to identify it:
@@
Parallel trends makes DiD distinct from causal designs that are based on statistical independence between treatment and potential outcomes. In designs like randomized trials or instrumental variables, the conditions—mean equalities across groups, for instance—that identify counterfactuals are often a statistical consequence of the randomness induced externally (Heckman2000). In contrast, parallel trends is just a convenient restriction on untreated potential outcome trends. It does not necessarily come from exogenous variation “outside the model.” In fact, because treatment adoption is often chosen by economic actors or policymakers “inside the model”, parallel trends need not hold automatically. For this reason, DiD analyses (correctly) devote significant attention to understanding whether parallel trends is plausible in a given application.
@@
We discuss how to generate empirical evidence about the plausibility of parallel trends below, but new theoretical findings about treatment choice behaviors or selection mechanisms also inform the plausibility of parallel trends. [^5] In fact, some of the earliest economic research on DiD methods examined exactly these questions (Ashenfelter1985; Heckman1985). These results explicitly connect DiD to behaviors, they provide grounding for stories about why a given DiD analysis is plausible (or not), and they discipline empirical tests of parallel trends. A full review of this literature is outside the scope of this paper, but some helpful broad themes emerge. [^6] For details on selection models and outcome models that are consistent with parallel trends see Ghanem2023; Marx2024; Chabeferret2015 applies these arguments to earnings dynamics. For instance, a trade-off exists between the information the agents who choose treatment know and how they act on it, and the time-series properties of untreated potential outcomes. At one extreme, consider someone who knows $Y_{i,t}(0)$ before and after treatment and can opt into or out of treatment based on it. Parallel trends can only hold in this case if, other than common shifts for every unit, $Y_{i,t}(0)$ is constant (Ghanem2023). In our Medicaid example, neither of these conditions—that state legislatures in 2013 knew their 2014 untreated mortality rates or that untreated mortality rates would all have shifted in parallel—is plausible. The opposite extreme is when treatment timing is random, in which case parallel trends holds without any time-series restrictions. Yet, in this case, DiD is not necessary, and more efficient estimators exist (Roth2023c). Against the political and economic backdrop of the early 2010s, the claim that state choices about expanding Medicaid were random is also implausible.
@@
Therefore, in realistic scenarios, parallel trends can only hold under some restrictions on the way untreated outcomes enter the treatment selection mechanism. As one example, imagine that treatment selection depends on the permanent component of $Y_{i,t}(0)$ (fixed effects) but not on shorter-term fluctuations (“shocks”). For instance, if state legislatures only knew and considered their long-run mortality levels when making their expansion decision, they would be following this kind of selection mechanism. Expansion and non-expansion states would then have large differences in the permanent part of untreated outcomes, which difference in equation (3.3), and so parallel trends would hold if shocks to $Y_{i,t}(0)$ had a stable mean. [^7] Some researchers may find easier to understand these as “parallel changes” rather than “parallel trends”. However, the use of “parallel trends” is now firmly established in the literature, and other influential work has used “changes-in-changes” to refer to an alternative estimator to classical DiD estimation (Athey2006). To avoid confusion, we use parallel trends throughout this paper. State legislatures, however, may have also known whether their 2013 mortality rates were especially high or low when considering expanding Medicaid. If the expansion choice is related to these 2013 mortality shocks as well as to fixed effect, parallel trends would hold only if one imposes stronger time-series restrictions on $Y_{i,t}(0)$. Ghanem2023 provides a fuller discussion of the selection/time-series trade-off and theory-driven templates to assess parallel trends, while Marx2024 discusses economic models that are and are not compatible with parallel trends. Another implication of the fact that DiD does not rely on statistical independence between $Y_{i,t}(0)$ and treatment status is that there is no guarantee that parallel trends holds across different transformations of $Y_{i,t}(0)$. As stated, it is simply an assumption about averages for a particular $Y_{i,t}(0)$. Roth2023b show that parallel trends is insensitive to functional form if and only if it holds between groups and across the distribution of $Y_{i,t}(0)$. This would entail assuming that either Medicaid adoption is random, the mortality distribution is constant between 2013 and 2014, or a mixture of the two cases. As these conditions are arguably ex-ante restrictive, our DiD analysis may depend on our choices to measure $Y_{i,t}$ in rates (deaths per 100,000) as opposed to logs, for example. One way to evaluate this measurement choice is to propose a theory that delivers it, though we recognize that this is not always possible. To assess whether cases where parallel trends holding for one functional form come at the cost of ruling other transformations, we recommend that researchers use the Roth2023b’s falsification tests for the null that parallel trends are insensitive to functional form. In our application, we do not reject the null that parallel trends is insensitive to functional form, with p-values above 0.80.
@@
The interplay between treatment selection and the properties of the outcome variable characterize the structural basis for a DiD analysis (see DiNardo2011) and engaging with them is essential to any DiD application. While every study will have its own institutions, choices, and outcomes to consider, a rigorous DiD analysis must provide a transparent discussion about the reliability of the underlying identification assumptions. If parallel trends is not plausible, one may be better off using an alternative research design.
3.3 Estimation and inference: 4 means or one regression?
@@
  • Notes: This table reports average county-level mortality rates (deaths among adults aged 20-64 per 100,000 adults) in 2013 (row 1) and 2014 (row 2) in states that expanded adult Medicaid eligibility in 2014 (columns 1 and 4) and states that have not expanded by 2019 (columns 2 and 5). The first three columns present unweighted averages and the second three columns present population-weighted averages. Columns 1, 2, 4, and 5 in the third row show time trends in mortality between 2013 and 2014 for each group of states. The first two rows of columns 3 and 6 show the cross-sectional gap in mortality between expansion and non-expansion states in 2013 and 2014. The entries in bold red text in row 3 show the simple 2 $\times$ 2 difference-in-differences estimates without weights (column 3) and with them (column 6)
The two across-time changes in equation (3.6) are in the third row of the table. Without weighting, average county-level mortality rates in expansion states rose by 9.3 deaths per 100,000 and 9.1 deaths in non-expansion states, so, after rounding, the DiD estimate of $ATT(2014)$ is 0.1 deaths per 100,000, implying that the average treatment effect of Medicaid expansion had on mortality in 2014 among counties that are part of an expansion state was an increase of 0.1 deaths per 100,000. In contrast, the DiD result using population weights suggests that Medicaid expansion caused a reduction of 2.6 deaths per 100,000 for the average adult in expansion states. [^8] Columns 3 and 6 show cross-group gaps in average mortality in each year. These can also be used to construct the DiD estimate by rearranging equation (3.6): $(\overline{Y}_{D=1,t=2014}-\overline{Y}_{D=0,t=2014})-(\overline{Y}_{D=1,t=2013}-\overline{Y}_{D=0,t=2013})$.
The same result can be obtained as the (weighted) least squares estimate of $\beta^{2\times 2}$ in the following linear regression specification (that only has data for $t=2013$ and $t=2014$):
@@
where $\beta$’s are unknown coefficients, $\varepsilon_{i,t}$ is an idiosyncratic term uncorrelated with $D_{i}$. To see why, let’s focus on the unweighted case ($\omega_{i}=1$), and write each of the four means in $\widehat{ATT}(2)$ in terms of the estimated coefficients from (3.7):
  • Sample average of $Y_{i,t}$ in post period for treatment group: $\overline{Y}_{D=1,t=2014}=\widehat{\beta_{0}}+\widehat{\beta_{1}}+\widehat{\beta_{2}}+\widehat{\beta}^{2\times 2}$.
  • Sample average of $Y_{i,t}$ in pre period for treatment group is $\overline{Y}_{D=1,t=2013}=\widehat{\beta_{0}}+\widehat{\beta_{1}}$.
  • Sample average of $Y_{i,t}$ in post period for comparison group is $\overline{Y}_{D=0,t=2014}=\widehat{\beta_{0}}+\widehat{\beta_{2}}$.
  • Sample average of $Y_{i,t}$ in pre period for comparison group is $\overline{Y}_{D=0,t=2013}=\widehat{\beta_{0}}$.
@@
Table 3 demonstrates this equivalence for both unweighted (column 1) and weighted (column 4) regressions. In fact, with balanced panel data, the estimate of $\beta^{2\times 2}$ is numerically the same if the regression instead contains fixed effects for each unit (columns 2 and 5) or if one regresses outcome changes on a constant and the treatment group dummy $D_{i}$ (columns 3 and 6). [^9] When population weights vary over time the equivalence between a by-hand DiD estimate and one that comes from a regression with unit fixed effects no longer holds.
@@
Unweighted Weighted Crude Mortality Rate $\Delta$ Crude Mortality Rate $\Delta$ (1) (2) (3) (4) (5) (6) Constant 474.0∗∗∗ 9.1∗∗∗ 376.4∗∗∗ 6.3∗∗∗ (4.3) (2.6) (7.6) (1.1) Medicaid Expansion -54.8∗∗∗ -53.7∗∗∗ (6.3) (11.5) Post 9.1∗∗∗ 6.3∗∗∗ (2.6) (1.1) Medicaid Expansion $\times$ Post 0.1 0.1 0.1 -2.6 -2.6 -2.6 (3.7) (3.7) (3.7) (1.5) (1.5) (1.5) County fixed effects No Yes No No Yes No Year fixed effects No Yes No No Yes No Notes: This table reports the regression 2 $\times$ 2 DiD estimate comparing counties that expand Medicaid in 2014 to counties that do not expand Medicaid by 2019, using only data for the years 2013 and 2014. Columns 1-3 report unweighted regression results, while columns 4-6 weight by county population aged 20-64 in 2013. Columns 1 and 4 report results from regressing the crude mortality rate for adults ages 20-64 on indicators for expansion states (Treat) and post-expansion year (Post), with the DiD estimate being the coefficient on the interaction term. Columns 2 and 5 report the corresponding results for the interaction term using county and year fixed effects. Finally, Columns 3 and 6 report the results of the long difference in county mortality rates on a treatment indicator. Standard errors (in parentheses) are clustered at the county level.
@@
The equivalence between calculating a $2\times 2$ DiD by hand or with a regression has appealing features. Regressions are simple to run, and they do the averaging and differencing behind the scenes. They also allow the use of statistical inference tools from ordinary least squares (OLS), which are themselves the subject of a large econometrics literature that is particularly important when it comes to clustering decisions (Wooldridge2003; Bertrand2004; Donald2007; Cameron2008; Conley2011; Abadie2020; Abadie2023).
@@
Many inference procedures exist for DiD-type analyses, arising from a combination of choices about the target parameter, details of the data structure and sampling process, and maintained assumptions about the structure of outcomes. In practice, one needs to determine and discuss the forms of uncertainty the standard errors are designed to capture; that is, what is (conceptually) being resampled and what may or may not vary across those resamples. As discussed in Abadie2020, these details come from the nature of the parameter of interest—whether the focus is on sample-specific average treatment effects or population-level average treatment effects—and the stochastic elements of the model that make the estimator random. Heuristically, this involves a thought experiment (or stochastic process) hypothesized to generate the random components of the model (or the data-generating process). Different inferential frameworks highlight different sources of uncertainty by resampling distinct model components and treating other components as fixed (non-random).
@@
Inferential frameworks on two extremes help cement these concepts. Design-based frameworks treat potential outcomes and covariates as non-random, focus on finite-population parameters (e.g., sample average treatment effects), and consider the allocation of treatment as the only source of the randomness in the model (imbensrubin2015). [^10] Traditionally, design-based inference procedures are justified when treatment assignment is fully random, which is a much stronger requirement than parallel trends. See Rambachan2022 for a discussion on design-based inference for quasi-experimental designs, including a discussion of the Medicaid expansion. The only thing that is random and thus varies across the hypothetical resamples from this point of view is the treatment allocation. On the other hand, a traditional sampling-based approach to inference presumes that we independently sample units from a superpopulation. In this case, it is customary to focus on population parameters (like $ATT(2)$), treat all variables in the model as random variables, and cluster standard errors at the level in which the (hypothesized) sampling was conducted. In this framework, every variable in the analysis—outcomes, covariates, and treatment—is randomly redrawn across the hypothetical resamples. A drawback of the sampling approach is that sometimes, it is unnatural to think of the data as a random sample from a well-defined population.
@@
A third popular approach to inference—the model-based approach—is more structural and involves taking a stand on the structure of the error component of the model (e.g., imposing a putative model for how shocks affect outcomes and their relationship with treatment and other variables in the model). The uncertainty reflected in this model-based setting entails a thought experiment in which different values of these shocks and the other random variables in the model are drawn from their joint distribution (Abadie2023). This model-based approach is common in econometrics, and it almost always takes the linear regression specification (or model, in this case) as the starting point of the analysis. Albeit this is often convenient, it is important to note that imposing model restrictions on the error component of the model necessarily imposes restrictions on treatment effect heterogeneity and on the relationship between potential outcomes; see Section 5 and Appendix A of Roth2023a for a discussion. Another challenge with the model-based approach is that it is hard to use this framework when adopting estimation strategies other than linear regressions, e.g., inverse probability weighting or doubly robust procedures that we will discuss in Session 4.4.
@@
Ultimately, as the discussion above highlights, each approach has pros and cons, and discussions about the best way to compute standard errors are complex and often ambiguous. As such, a detailed treatment of the topic is outside the scope of this paper. We emphasize that such a discussion is intrinsically context-specific, requiring information about the sampling process, the research design and target parameters, what is treated as fixed and random, and the structure of the error components of their models, among other factors. We refer interested readers to Abadie2020; Abadie2023 and Section 5 of Roth2023a for discussions on these topics, though we also emphasize that further methodological research in this area is warranted. For the remainder of this article, we adopt a sampling perspective for uncertainty and cluster our standard errors at the county level. In our context, this is compatible with treating all variables as random, including treatment groups and potential outcomes. It also allows us to avoid (a) making time-series dependence restrictions on potential (and realized) outcomes—as we are in a short-panel framework with a large number of units and a fixed number of time periods—and (b) taking an explicit strand on the structure of error components of the model, which is particularly appealing as the starting point of our analysis are potential outcomes and not regression models. It is also worth mentioning that as our treatment in the empirical example is assigned at the state level, clustering at the county level would also be compatible with treating state-specific shocks as fixed (or conditioned on) and assessing if they lead to violations of parallel trends (Roth2023a, Section 5.1). Clustering at the state level would be justified using a design-based perspective (Rambachan2022), though that would require us to treat potential outcomes as fixed (which we do not in this paper). Our choice of inference procedure is not without controversies, and other inferential approaches may also be rationalized under additional auxiliary structures. However, we do not follow that path in this paper.
@@
We conclude this section by stressing that the appeal of using regressions like (3.7) to estimate ATT in DiD designs comes from the fact that it is numerically equivalent to the “by-hand” DiD estimator (3.6), which was explicitly derived from the $ATT(t)$ and the parallel trends assumption. This ensures that the regression specification respects the underlying identifying assumptions and estimates the desired target parameter. Unfortunately, the tight connection between (TWFE) regressions and DiD designs breaks under more complex setups that are ubiquitous in practice. We now turn to some of these issues and how approaching them from the point of view of $2\times 2$ building blocks can guide good econometric practices.

### v2

@@ 3.2 Identifying assumptions: Parallel trends
A research design is a strategy—a set of assumptions—to identify and estimate specific target parameters. Many different assumptions can identify the missing counterfactual for $ATT(2014)$ in the Medicaid example. For example, mean independence between $Y_{i,2014}(0)$ and $D_{i}$ implies that the counterfactual equals average 2014 mortality rates in non-expansion counties ($\mathop{}\!\mathbb{E}_{\omega}[Y_{i,2014}(0)|D_{i}=0]$). Under this assumption, which essentially entails assuming that Medicaid expansion is as good as random, the cross-sectional mortality gap in 2014 between expansion and non-expansion counties is the $ATT(2014)$. Similarly, time invariance of $Y_{i,t}(0)$ among expansion counties (plus the fact that we ruled out anticipatory behavior) implies that the counterfactual equals 2013 mortality rates in expansion counties ($\mathop{}\!\mathbb{E}_{\omega}[Y_{i,2013}(0)|D_{i}=1]$). Under this assumption, which essentially rules out non-treatment-related changes in the outcome variable, the “time trend” in average mortality in expansion counties is the $ATT(2014)$.
DiD comes from an alternative assumption that identifies the relevant counterfactual even when the average untreated potential outcome differs across treatment groups (which violates mean independence) and changes over time (which violates time invariance). The parallel trends assumption states that in the absence of treatment, the average outcome evolution is the same among treated and comparison groups. For general assumptions and results, we denote time periods by $t=1,2$, but continue to be explicit about which years are being used when we reference the Medicaid example.
@@
The (weighted) average change of $Y_{i,t=2}(0)$ from $Y_{i,t=1}(0)$ is the same between treated and comparison groups; that is,
@@
If the parallel trends assumption holds, then it is easy to construct $\mathop{}\!\mathbb{E}_{\omega}[Y_{i,t=2}(0)|D_{i}=1]$ from observable quantities—that is, to identify it:
@@
Parallel trends makes DiD distinct from causal designs that are based on statistical independence between treatment and potential outcomes. In designs like randomized trials or instrumental variables, the conditions—mean equalities across groups, for instance—that identify counterfactuals are often a statistical consequence of the randomness induced externally (Heckman2000). In contrast, parallel trends is just a restriction on untreated potential outcome trends. It does not necessarily come from exogenous variation “outside the model.” In fact, because treatment adoption is often chosen by economic actors or policymakers “inside the model,” parallel trends need not hold. For this reason, DiD analyses (correctly) devote significant attention to evaluating parallel trends.
@@
We discuss how to generate empirical evidence about the plausibility of parallel trends below, but new theoretical findings about treatment choice behaviors or selection mechanisms also inform the plausibility of parallel trends. [^4] In fact, some of the earliest economic research on DiD methods examined exactly these questions (Ashenfelter1985; Heckman1985). These results explicitly connect DiD to behaviors, they provide grounding for stories about why a given DiD analysis is internally valid (or not), and they discipline empirical tests of parallel trends. A full review of this literature is outside the scope of this paper, but some helpful broad themes emerge. [^5] For details on selection models and outcome models that are consistent with parallel trends, see Ghanem2023 and Marx2024; Chabeferret2015 applies these arguments to earnings dynamics. For instance, a trade-off exists between the information the agents who choose treatment know and how they act on it, and the time-series properties of untreated potential outcomes. At one extreme, consider someone who knows $Y_{i,t}(0)$ before and after treatment and can opt into or out of treatment based on it. Parallel trends can only hold in this case if, other than common shifts for every unit, $Y_{i,t}(0)$ is constant (Ghanem2023). In our Medicaid example, neither of these conditions—that state legislatures in 2013 knew their 2014 untreated mortality rates or that untreated mortality rates would all have shifted in parallel—is plausible. The opposite extreme is when treatment timing is random, in which case parallel trends holds without any time-series restrictions. Yet, in this case, DiD is not necessary, and more efficient estimators exist (Roth2023c). Against the political and economic backdrop of the early 2010s, it is clear that Medicaid expansion decisions were not random.
@@
Therefore, in realistic scenarios, parallel trends only holds if some restrictions on the way untreated outcomes enter the treatment selection mechanism hold as well. As one example, imagine that treatment selection depends on the permanent component of $Y_{i,t}(0)$ (fixed effects) but not on shorter-term fluctuations (“shocks”). For instance, if state legislatures knew and considered their long-run mortality levels only when making their expansion decision, they would be following this kind of selection mechanism. Expansion and non-expansion states might then have large differences in the permanent part of untreated outcomes, which cancel in equation (3.3), and so parallel trends would hold if shocks to $Y_{i,t}(0)$ had a time invariant mean conditional on the fixed effects. [^6] Some researchers may find easier to understand these as “parallel changes” rather than “parallel trends.” However, the use of “parallel trends” is now firmly established in the literature, and other influential work has used “changes-in-changes” to refer to an alternative estimator to classical DiD estimation (Athey2006). To avoid confusion, we use “parallel trends” throughout this paper. State legislatures, however, may have also known whether their 2013 mortality rates were especially high or low when considering expanding Medicaid. If the expansion choice is related to these 2013 mortality shocks as well, parallel trends would hold only if stronger time-series restrictions on $Y_{i,t}(0)$ hold. Ghanem2023 provide a fuller discussion of the selection/time-series trade-off and theory-driven templates to assess parallel trends, while Marx2024 discuss economic models that are and are not compatible with parallel trends.
Another implication of the fact that DiD does not rely on statistical independence between $Y_{i,t}(0)$ and treatment status is that there is no guarantee that parallel trends holds across different transformations of $Y_{i,t}(0)$. As stated, it is simply an assumption about averages for a particular $Y_{i,t}(0)$. Roth2023b show that parallel trends is insensitive to functional form if and only if it holds between groups and across the distribution of $Y_{i,t}(0)$. This can only be true if Medicaid adoption is random, the mortality distribution is constant between 2013 and 2014, or a combination of the two cases. As these conditions are arguably ex-ante implausible, our DiD analysis may depend on our choices to measure $Y_{i,t}$ in rates (deaths per 100,000) as opposed to logs, for example. One approach to this measurement choice is to propose and evaluate a theory that delivers it, though we recognize that this is not always possible. To assess whether cases where parallel trends holding for one functional form come at the cost of ruling other transformations, we recommend that researchers use the Roth2023b’s (Roth2023b) falsification tests for the null that parallel trends are insensitive to functional form. In our application, we do not reject the null that parallel trends is insensitive to functional form, with p-values above 0.80.
@@
The interplay between treatment selection and the properties of the outcome variable characterize the structural basis for a DiD analysis (see DiNardo2011) and engaging with them is essential to any DiD application. While every study will have its own institutions, choices, and outcomes to consider, a rigorous DiD analysis must provide a transparent discussion about the reliability of the underlying identification assumptions. If parallel trends requires implausible behavioral restrictions, one may be better off using an alternative research design.
3.3 Estimation and inference: Four means or one regression?
@@
  • This table reports average county-level mortality rates (deaths among adults aged 20-64 per 100,000 adults) in 2013 (top row) and 2014 (middle row) in states that expanded adult Medicaid eligibility in 2014 (columns 1 and 4, 978 counties) and states that have not expanded by 2019 (columns 2 and 5, 1,222 counties). The first three columns present unweighted averages and the second three columns present population-weighted averages. Columns 1, 2, 4, and 5 in the third row show time trends in mortality between 2013 and 2014 for each group of states. The first two rows of columns 3 and 6 show the cross-sectional gap in mortality between expansion and non-expansion states in 2013 and 2014. The entries in red text in the bottom row show the simple 2 $\times$ 2 difference-in-differences estimates without weights (column 3) and with them (column 6)
The two across-time changes in equation (3.6) are in the third row of the table. Without weighting, average county-level mortality rates in expansion states rose by 9.3 deaths per 100,000 and 9.1 deaths in non-expansion states, so after rounding, the DiD estimate of $ATT(2014)$ is 0.1 deaths per 100,000. This result implies that the average treatment effect of Medicaid expansion on mortality in 2014 among counties that are part of an expansion state was an increase of 0.1 deaths per 100,000. In contrast, the DiD result using population weights suggests that Medicaid expansion caused a reduction of 2.6 deaths per 100,000 for the average adult in expansion states. [^7] Columns 3 and 6 show cross-group gaps in average mortality in each year. These can also be used to construct the DiD estimate by rearranging equation (3.6): $(\overline{Y}_{D=1,t=2014}-\overline{Y}_{D=0,t=2014})-(\overline{Y}_{D=1,t=2013}-\overline{Y}_{D=0,t=2013})$.
The same result can be obtained as the (weighted) least squares estimate of $\beta^{2\times 2}$ in the following linear regression specification (which only uses data from $t=2013$ and $t=2014$):
@@
where $\beta$’s are unknown coefficients and $\varepsilon_{i,t}$ is an idiosyncratic term uncorrelated with $D_{i}$. To see why, let’s focus on the unweighted case ($\omega_{i}=1$) and write each of the four means in $\widehat{ATT}(2)$ in terms of the estimated coefficients from (3.7):
  • Sample average of $Y_{i,t}$ in post-period for treatment group is $\overline{Y}_{D=1,t=2014}=\widehat{\beta_{0}}+\widehat{\beta_{1}}+\widehat{\beta_{2}}+\widehat{\beta}^{2\times 2}$.
  • Sample average of $Y_{i,t}$ in pre-period for treatment group is $\overline{Y}_{D=1,t=2013}=\widehat{\beta_{0}}+\widehat{\beta_{1}}$.
  • Sample average of $Y_{i,t}$ in post-period for comparison group is $\overline{Y}_{D=0,t=2014}=\widehat{\beta_{0}}+\widehat{\beta_{2}}$.
  • Sample average of $Y_{i,t}$ in pre-period for comparison group is $\overline{Y}_{D=0,t=2013}=\widehat{\beta_{0}}$.
@@
Table 3 demonstrates this equivalence for both unweighted (column 1) and weighted (column 4) regressions. In fact, with balanced panel data, the estimate of $\beta^{2\times 2}$ is numerically the same if the regression instead contains fixed effects for each unit (columns 2 and 5) or if one regresses outcome changes on a constant and the treatment group dummy $D_{i}$ (columns 3 and 6). [^8] When population weights vary over time, the equivalence between a by-hand DiD estimate and one that comes from a regression with unit fixed effects no longer holds. Provided that one adopts the same notion of uncertainty, standard errors also coincide; such as when one clusters the standard errors at the county level.
@@
Unweighted Weighted Crude Mortality Rate $\Delta$ Crude Mortality Rate $\Delta$ (1) (2) (3) (4) (5) (6) Constant 474.0∗∗∗ 9.1∗∗∗ 376.4∗∗∗ 6.3∗∗∗ (4.3) (2.6) (7.6) (1.1) Medicaid Expansion -54.8∗∗∗ -53.7∗∗∗ (6.3) (11.5) Post 9.1∗∗∗ 6.3∗∗∗ (2.6) (1.1) Medicaid Expansion $\times$ Post 0.1 0.1 0.1 -2.6 -2.6 -2.6 (3.7) (3.7) (3.7) (1.5) (1.5) (1.5) County fixed effects No Yes No No Yes No Year fixed effects No Yes No No Yes No This table reports the regression 2 $\times$ 2 DiD estimate comparing 978 counties in states that expanded Medicaid in 2014 with 1,222 counties in states that did not expand Medicaid by 2019, using only data for the years 2013 and 2014. Columns 1-3 report unweighted regression results, while columns 4-6 weight by county population aged 20-64 in 2013. Columns 1 and 4 report results from regressing the crude mortality rate for adults ages 20-64 on indicators for expansion states (Treat) and post-expansion year (Post); the DiD estimate is the coefficient on the interaction term. Columns 2 and 5 report the corresponding results for the interaction term using county and year fixed effects. Finally, Columns 3 and 6 report the results of the long difference in county mortality rates on a treatment indicator. Standard errors (in parentheses) are clustered at the county level.
@@
The equivalence between calculating a $2\times 2$ DiD by hand or with a regression has appealing features. Regressions are simple to run, and they do the averaging and differencing behind the scenes. They also allow the use of statistical inference tools from ordinary least squares (OLS), which are themselves the subject of a large econometrics literature that is particularly important when it comes to standard error estimation (Wooldridge2003; Bertrand2004; Donald2007; Cameron2008; Conley2011; Abadie2020; Abadie2023).
@@
Many inference procedures exist for DiD-type analyses, arising from a combination of choices about the target parameter, details of the data structure and sampling process, and maintained assumptions about the structure of outcomes. In practice, one needs to determine and discuss the forms of uncertainty the standard errors are designed to capture—that is, what is (conceptually) being resampled and what may or may not vary across those resamples. As discussed in Abadie2020, these details come from the nature of the parameter of interest—whether the focus is on sample-specific average treatment effects or population-level average treatment effects—and the stochastic elements of the model that make the estimator random. Heuristically, this involves a thought experiment (or stochastic process) hypothesized to generate the random components of the model (or the data-generating process). Different inferential frameworks highlight different sources of uncertainty by resampling distinct model components and treating other components as fixed (non-random).
@@
Inferential frameworks on two extremes help cement these concepts. Design-based frameworks treat potential outcomes and covariates as non-random, focus on finite-population parameters (e.g., sample average treatment effects), and consider the allocation of treatment as the only source of the randomness in the model (imbensrubin2015). [^9] Traditionally, design-based inference procedures are justified when treatment assignment is fully random, which is a much stronger requirement than parallel trends. See Rambachan2022 for a discussion on design-based inference for quasi-experimental designs, including a discussion of the Medicaid expansion. The only thing that is random and thus varies across the hypothetical resamples from this point of view is the treatment allocation. On the other hand, a traditional sampling-based approach to inference presumes that we independently sample units from a superpopulation. In this case, it is customary to focus on population parameters (like $ATT(2)$), treat all variables in the model as random variables, and compute standard errors by clustering at the level in which the (hypothesized) sampling was conducted. In this framework, every variable in the analysis—outcomes, covariates, and treatment—is randomly redrawn across the hypothetical resamples. A drawback of the sampling approach is that sometimes, it is unnatural to think of the data as a random sample from a well-defined population.
@@
A third popular approach to inference—the model-based approach—is more structural and involves taking a stand on the structure of the error component of the model (e.g., imposing a putative model for how shocks affect outcomes and their relationship with treatment and other variables in the model). The uncertainty reflected in this model-based setting entails a thought experiment in which different values of these shocks and the other random variables in the model are drawn from their joint distribution (Abadie2023). This model-based approach is common in econometrics, and it almost always takes the linear regression specification (or model, in this case) as the starting point of the analysis. Although this is often convenient, it is important to note that imposing model restrictions on the error component of the model necessarily imposes restrictions on treatment effect heterogeneity and on the relationship between potential outcomes; see Section 5 and Appendix A of Roth2023a for a discussion. Another challenge with the model-based approach is that it is hard to use this framework when adopting estimation strategies other than linear regressions—for example, inverse probability weighting or doubly robust procedures—which we will discuss in Session 4.4.
@@
Ultimately, as the discussion above highlights, each approach has pros and cons, and discussions about the best way to compute standard errors are complex and intrinsically context-specific. A detailed treatment of the topic is outside the scope of this paper. It requires information about the sampling process, the research design and target parameters, what is treated as fixed and random, and the structure of the error components of their models (e.g., presence of spatial or serial correlation), among other factors. We refer interested readers to Abadie2020; Abadie2023 and Section 5 of Roth2023a for discussions on these topics, though we also emphasize that further methodological research in this area is warranted. For the remainder of this article, we adopt a sampling perspective for uncertainty and cluster our standard errors at the county level. In our context, this is compatible with treating all variables as random, including treatment groups and potential outcomes. It also allows us to avoid (a) making time-series dependence restrictions on potential (and realized) outcomes—as we are in a short-panel framework with a large number of units and a fixed number of time periods—and (b) taking an explicit stand on the structure of error components of the model, which is particularly appealing as the starting point of our analysis is potential outcomes rather than regression models. It is also worth mentioning that as our treatment in the empirical example is assigned at the state level, clustering at the county level would also be compatible with treating state-specific shocks as fixed (or conditioned on) and assessing if they lead to violations of parallel trends (Roth2023a, Section 5.1). Clustering at the state level would be justified if we were using a design-based perspective (Rambachan2022), though that would require us to treat potential outcomes as fixed (which we do not do in this paper). Our choice of inference procedure is not without controversies, and other inferential approaches may also be rationalized.
@@
We conclude this section by stressing that the appeal of using regressions like (3.7) to estimate ATT in DiD designs comes from the fact that their numerical equivalence to the “by-hand” DiD estimator (3.6), which was explicitly derived from the $ATT(t)$ and the parallel trends assumption. This ensures that the regression specification respects the underlying identifying assumptions and estimates the desired target parameter. Unfortunately, the tight connection between (TWFE) regressions and DiD designs breaks under more complex setups that are ubiquitous in practice. We now turn to some of these issues and how approaching them from the point of view of $2\times 2$ building blocks can guide good econometric practices.